OpenStudy (anonymous):

(x^2+3x)/(x^2-4) partial fraction with steps please

OpenStudy (kc_kennylau):

\[\frac{x^2+3x}{x^2-4}\equiv\frac{A}{x-2}+\frac{B}{x+2}\]

OpenStudy (anonymous):

I keep getting 1-5/2(x+2)-1/2(x-2)

OpenStudy (kc_kennylau):

wait, i mean: \[\frac{x^2+3x}{x^2-4}\equiv x\left(\frac A{x-2}+\frac B{x+2}\right)\]

OpenStudy (anonymous):

I don't know what I'm doing wrong

OpenStudy (kc_kennylau):

Ax^2+Bx^2=x^2 2Ax-2Bx=3x

OpenStudy (kc_kennylau):

A+B=1 2A-2B=3

OpenStudy (kc_kennylau):

2(1)-(2): 4B=-1, B=-0.25

OpenStudy (kc_kennylau):

(1): A=1.25

OpenStudy (kc_kennylau):

\[\frac{x^2+3x}{x^2-4}\equiv\frac{1.25x}{x-2}-\frac{0.25x}{x+2}\]

OpenStudy (kc_kennylau):

what did you get? please use proper brackets

OpenStudy (anonymous):

The answer says 1+(5/2)(x-2)+1/2(x+2) but I get 1-(5/2)(x+2)-1/2(x-2)

OpenStudy (kc_kennylau):

wait, can you use a divide sign to separate the numerator and the denominator

OpenStudy (kc_kennylau):

i can hardly interpret what you wrote

OpenStudy (anonymous):

5/2 is a fraction as well as 1/2

OpenStudy (rsadhvika):

you need to divide first to make denominator's degree less than numerator

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

Hence the 1

OpenStudy (kc_kennylau):

Do you know what you wrote \[1+\frac52(x-2)+\frac12(x+2)\]

OpenStudy (rsadhvika):

\(\large \frac{x^2+3x}{x^2-4} = \frac{x^2-4 + 4 + 3x}{x^2-4} = 1 + \frac{3x+4}{x^2-4}\)

OpenStudy (anonymous):

Yes I got to that part

OpenStudy (anonymous):

Wait almost

OpenStudy (rsadhvika):

\(\large \frac{3x+4}{x^2-4} = \frac{A}{x-2} + \frac{B}{x+2}\) use cover up and find A, B

OpenStudy (anonymous):

I think the signs are different

OpenStudy (anonymous):

That's my mistake

OpenStudy (rsadhvika):

The answer says 1+(5/2)(x-2)+1/2(x+2) but I get 1-(5/2)(x+2)-1/2(x-2) ^^^^^ ^^^^^

OpenStudy (rsadhvika):

work it again, not just signs.. values above A and B also flipped

OpenStudy (anonymous):

Okay I believe I've done the division in the first step in reverse Thank you

OpenStudy (rsadhvika):

cool ^_^

OpenStudy (anonymous):

Now I can integrate

OpenStudy (rsadhvika):

yes, that should be trivial, just the ln 's

OpenStudy (rsadhvika):

:)